# Markarian 501 reincarnates with vigor as a temporary EHBL during VHE flaring in July 2014

### v1

@@ Abstract
Markarian 501, a well known high energy BL Lac object, has exhibited several epochs of very high energy (VHE) gamma-ray flaring events when its synchrotron peak frequency shifted above $10^{17}$ Hz, a signature of extreme behavior. From July 16 to July 31, 2014 such flaring events were observed for 15 days by various telescopes. On July 19 (MJD 56857.98), the X-ray outburst from the source was at its highest and on the same day an intriguing narrow peak-like feature around 3 TeV was observed by MAGIC telescopes, a feature inconsistent with standard interpretations. Using the well known two-zone photohadronic model, we study these VHE gamma-ray spectra on a day-by-day basis and offer explanation. Our two-zone photohadronic scenario shows that, on MJD 56857.98, the peak-like feature appears at a cutoff energy of $E^{c}_{\gamma}=3.18$ TeV. Below this energy the observed VHE spectrum increases slowly and above $E^{c}_{\gamma}$ it falls faster, thus resulting in a mild peak-like feature.
High energy astrophysics (739), Blazars (164), Gamma-rays (637), Relativistic jets (1390), BL Lacertae objects (158)
[…]
@@ 1 Introduction
Markarian 501 (Mrk 501) is a high synchrotron peaked BL Lac object (HBL) at a redshift of $z=0.034$ (Ulrich et al., 1975). It is the second well known bright HBL after Markarian 421 (Mrk 421) and is one of the brightest extragalactic sources in the X-ray/TeV sky. Since its discovery in VHE by Whipple telescopes in 1996 (Abdo et al., 2009), several major outburst in multi-TeV have been observed and has been the target of many multiwavelength campaigns mainly covering VHE flaring activities (Kataoka et al., 1999; Albert et al., 2008; Kranich, 2009; Aleksić et al., 2015). Mrk 501 is one of the few blazars which can be monitored even during the low emission periods. To understand different emission mechanisms and their characterizations, it is essential to have long-term and multiwavelength observations of the spectral energy distribution (SED) of the blazars. Also, simultaneous correlation studies in different wavelengths of the SED during the flaring can help us to constrain the emission mechanisms. Motivated by these goals, several multi-wavelength and multi-instrument observation campaigns were organized. In 1997, from March to October, Mrk 501 had an unprecedented flare in VHE gamma-rays and the integral flux reached up to four times the magnitude of the Crab Nebula (CU) flux (Petry et al., 2000). Also, it was extremely active in the X-ray region during this period and the synchrotron peak had a dramatic shift to energies above 100 keV, which is more than two orders of magnitude compared to that of the typical non-flaring state.
@@
During May to July 2005, Major Atmospheric Gamma-Ray Imaging Cherenkov (MAGIC) telescopes made observations for 30 nights and flaring above 100 GeV was observed with an order of magnitude flux variation (Albert et al., 2007). On the nights of June 30 and July 9, flux-doubling in VHE band was observed within a time span of about 2 minutes. This is the fastest flux variation ever observed from Mrk 501 since its discovery. During this observation period, the synchrotron peak and the VHE peak were shifted towards higher energy limits. As a part of the long term multiwavelength campaign, Mrk 501 was observed by $\sim 30$ different instruments and covered the entire electromagnetic spectrum (Aliu et al., 2016; Ahnen et al., 2017) from March 15 to August 1, 2009.
@@
Between March and July of 2012, another multiwavelength campaign was undertaken when more than 25 different telescopes, including MAGIC, and Very Energetic Radiation Imaging Telescope Array System (VERITAS), took part and fluxes above 0.2 TeV were observed (Ahnen et al., 2018). The highest peak flux was recorded on June 9, and most importantly, the X-ray spectrum and the VHE gamma-ray spectrum were observed to be extremely hard. During this period the synchrotron peak was above 5 keV and the Synchrotron Self-Compton (SSC) peak was above 0.5 TeV. The TeV Atmospheric Cherenkov Telescope with Imaging Camera (TACTIC) at Mt Abu, Rajastan, India, observed Mrk 501 from 15 April to 30 May, 2012 at energies above 850 GeV energy for 70.6 h. Between 22 to 27 May, 2012, TACTIC telescope observed relatively high VHE flux in the energy range between 850 GeV and 17.24 TeV (Chandra et al., 2017).
@@
Mrk 501 was monitored in the year 2014 with the First G-APD Cherenkov Telescope (FACT) (Anderhub et al., 2013; Biland et al., 2014; Dorner et al., 2015). Several high flaring states were observed. This triggered the HESS telescopes to undertake further observations. Enhanced VHE gamma-ray flux, comparable to the exceptional flare of 1997, was observed on the night of June 23-24, 2014 (MJD 56832) with the HESS phase-II telescopes. The energy of the VHE photons extended up to 20 TeV and the flux had a rapid variability. Estimates indicate that this is the highest flux ever observed from Mrk 501 with the HESS telescopes (Abdalla et al., 2019).
@@
Again, in July of 2014, flaring in X-ray and VHE gamma-ray were observed for about two weeks (from 16th to 31st) by the MAGIC telescopes (MAGIC Collaboration et al., 2020). The X-ray flaring was found to be exceptionally high, similar to the historic 1997 flaring. This amounts to the largest flux ever detected so far with the X-ray Telescope (XRT) on board the Neil Gehrels Swift Gamma-ray Burst Observatory (Gehrels et al., 2004) in its operational history. Also, the VHE gamma-ray emission above 0.15 TeV during this time was high, about 0.5 to 2 times of the Crab nebula flux. On July 19 (MJD 56857.98), the day of the highest X-ray activity, a narrow peak-like feature at $\sim 3$ TeV was observed by MAGIC telescopes (MAGIC Collaboration et al., 2020).
@@
In the Universe, the most dominant sources of extragalactic $\gamma$-rays are the blazars (Romero et al., 2017). Their SED is characterized by two non-thermal peaks (Abdo et al., 2010), the synchroton peak is in the infrared to X-ray frequency range and the SSC peak is in the X-ray to gamma-ray range. The BL Lac objects belong to the blazar family and are classified according to the position of their synchroton peak frequency $\nu^{p}_{s}$ in the SED. The classification list is as, low-energy peaked blazars (LBLs, $\nu^{p}_{s}<10^{14}\,\mathrm{Hz}$), intermediate-energy peaked blazars (IBLs, $10^{14}\,\mathrm{Hz}<\nu^{p}_{s}<10^{15}\,\mathrm{Hz}$), high-energy peaked blazars (HBLs, $10^{15}\,\mathrm{Hz}<\nu^{p}_{s}<10^{17}\,\mathrm{Hz}$ (Padovani & Giommi, 1995; Abdo et al., 2010; Boettcher et al., 2013)) and extreme high-energy peaked blazars (EHBLs, $\nu^{p}_{s}>10^{17}\,\mathrm{Hz}$ (Costamante et al., 2001)) with extreme spectral properties.
@@
Low luminosity and limited variability of EHBLs hinder the detection of their gamma-rays by the current generation Imaging Atmospheric Cerenkov Telescopes (IACTs). Thus, only a handful number of EHBLs have been detected so far and well known among them are 1ES 0229+200, 1ES 0347-232, RGB J0710+591 and 1ES 1101-232 (Costamante et al., 2018). It is observed that, at least three nearby extensively studied HBLs, Mrk 421, Mrk 501 and 1ES 1959+650 have shown temporary EHBL (tEHBL) like behavior during several flaring epochs  (Sahu et al., 2020, 2021a, 2021b). During tEHBL-like behavior, the synchrotron peaks have shifted above $10^{17}$ Hz and the TeV spectra are harder. Also, due to the complex behavior of these sources with different spectral properties at VHE gamma-rays, the possibility of different sub-classes within the EHBL class can not be ruled out (Foffano et al., 2019).
@@
With the standard one-zone leptonic SSC model, it is dificult to explain the VHE spectrum of EHBL, since the leptonic model predicts softer SSC spectrum in the Klein-Nishina regime, and is contrary to observations (Paggi et al., 2009). However, one can circumvent this problem by adopting unrealistic parameters, such as, large bulk Lorentz factor, large electron Lorentz factor and a small magnetic field (Ahnen et al., 2018; Acciari et al., 2020). Alternative solutions are also proposed, such as: the two-zone leptonic model, Inverse Compton scattering of the electrons with the cosmic microwave background, the spine-layer structured jet model and different hadronic models (Böttcher et al., 2008; Acciari et al., 2019; Ahnen et al., 2018).
@@
VHE gamma-rays from the extragalactic sources get attenuated by interacting with the extragalactic background light (EBL) and this attenuation is energy dependent. The shape of the observed spectrum changes at very high energies, so, EBL corrections to the observed spectrum have to be taken care of (Franceschini et al., 2008; Dominguez et al., 2011; Tavecchio, 2014; Tavecchio & Bonnoli, 2015). EHBLs have extreme spectral properties in TeV energy range, and probably these are ideal probes for EBL.
The multiwavelength SED of the HBL, Mrk 501, during several flaring periods are explained well using the leptonic models and hardonic models(Böttcher et al., 2008; Acciari et al., 2019; Ahnen et al., 2018; Sahu, 2019; Sahu et al., 2020). However, Mrk 501 had VHE flaring during 2005 and 2012 when the synchrotron peak had shifted above 1 keV, clearly a signature of EHBL-like behavior. In the context of leptonic model, explaining such shift and simultaneously explaining the VHE flaring event is problematic. We have used the two-zone photohadronic model to explain the tEHBL-like bahavior of these flaring events very well. We show that the photon background in the low energy tail region of the SSC spectrum is not a single power-law as in the case of HBLs, but consisists of two power-laws, separating it in to two different zones. Similarly, the extensively studied nearby HBLs, Mrk 421 and 1ES 1959+650 have also shown tEHBL-like behavior occasionally, and using the two-zone photohadronic model, we have successfully explained their VHE flaring events. In this work our goal is to explain the VHE flaring events observed during July 16 to July 31, 2014, and particularly, to interpret the unusual peak-like feature at about 3 TeV observed on July 19 (MJD 56857.98) in the context of the two-zone photohadronic model.
@@
The paper is organized as follows. In section 2, the two-zone photohadronic model is briefly discussed. In section 3, we discuss the flaring events of Mrk 501 during July 16 to July 31, 2014 (MJD 56854–56869). A detail analysis of the VHE events observed by MAGIC for about two weeks are discussed in section 4 and a brief discussion is given in section 5.
@@ 2 Two-Zone Photohadronic Model
The photohadronic model is very successful in explaining the VHE spectra from several HBLs (Sahu et al., 2013; Sahu, 2019; Sahu et al., 2019). This model is based on the fact that the low and high energy peaks of the blazar are produced from the synchrotron process of the electrons and inverse Compton scattering of electrons in the jet environment respectively (the leptonic model). The region where these photons are emitted is a blob of comoving radius $R^{\prime}_{b}$ (where implies comoving frame), with a bulk Lorentz factor $\Gamma$ and a Doppler factor $\mathcal{D}$. For HBLs, $\Gamma\simeq\mathcal{D}$. In the photohadronic scenario, accelerated protons are injected in the jet with a differential spectrum $dN/dE_{p}\propto E^{-\alpha}_{p}$, where the spectral index $\alpha\geq 2$ (Dermer & Schlickeiser, 1993). These protons interact with the background seed photons to produce $\Delta$-resonances, $p\gamma\rightarrow\Delta^{+}$. Decay of the $\Delta^{+}$ to neutral pion produces observed gamma-rays. The kinematical condition to produce $\Delta^{+}$ is given by,
$E_{p}\epsilon_{\gamma}=\frac{0.32\ \Gamma^{2}}{(1+z)^{2}}\ \mathrm{GeV^{2}},$ (1)
where $E_{p}$ and $\epsilon_{\gamma}$ are respectively the observed proton energy, the seed photon energy and $z$ is the redshift of the source. The energy of the observed gamma-ray $E_{\gamma}$ from the $\pi^{0}$ decay carries 10% of the original proton energy.
It is observed that in the standard jet scenario of the BL lac objects, the $p\gamma\rightarrow\Delta^{+}$ process is inefficient due to the low background photon density (Pjanka et al., 2017). To void this problem in the photohdronic model we assume that during the VHE flaring of the source, an internal compact jet is formed within the larger jet (Sahu et al., 2016, 2018). The inner jet has radius $R^{\prime}_{f}$ and the outer jet has radius $R^{\prime}_{b}$ and $R^{\prime}_{f}<R^{\prime}_{b}$. Also, the photon density in the inner jet region, $n^{\prime}_{\gamma,f}$, is larger than the photon density in the outer jet region, $n^{\prime}_{\gamma}$ ($n^{\prime}_{\gamma,f}\,>\,n^{\prime}_{\gamma}$). During the VHE flaring, the accelerated high energy protons interact with the seed photons in inner jet region to produce $\Delta$-resonance. Due to high density photons in the inner jet region, the efficiency of the production of $\Delta$-resonance is relatively high as compared to the standard jet scenario and this avoids the necessity for super-Eddington luminosity of protons.
The inner jet region is hidden and the photon density $n^{\prime}_{\gamma,f}$ is unknown. But the photon density in the outer jet region is known. Also, due to the adiabatic expansion of the inner jet, its photon density decreases when it crosses into the outer jet. To connect the inner and the outer region photon densities we assume a scaling relation given by (Sahu et al., 2019)
@@
From the previous analysis of the VHE spectra of dozens of HBLs in the photohadronic model, we observe that to produce the observed $E_{\gamma}$, the background seed photon energy $\epsilon_{\gamma}$ always lies in the low energy tail region of the SSC spectrum and the SSC flux is a perfect power-law, given by $\Phi_{SSC}\propto\epsilon^{\beta}_{\gamma}\propto E^{-\beta}_{\gamma}$, with $\beta$ the photon spectral index (Sahu, 2019). Thus the observed GeV-TeV photon flux from a HBL is given as
$F_{\gamma,obs}(E_{\gamma})=F_{0}\left(\frac{E_{\gamma}}{TeV}\right)^{-\delta+3}\,e^{-\tau_{\gamma\gamma}(E_{\gamma},z)}=F_{\gamma,in}(E_{\gamma})\,e^{-\tau_{\gamma\gamma}(E_{\gamma},z)}.$ (3)
@@
In the above equation, $F_{0}$ is the flux normalization, $F_{\gamma,in}$ is the intrinsic VHE gamma-ray flux and the exponential factor corresponds to the survival probability of VHE photon due to its interaction with the EBL photons which has the optical depth $\tau_{\gamma\gamma}$. For the EBL correction to the VHE spectrum, here, we use the EBL model of Franceschini et al. (2008). The exponent $\delta=\alpha+\beta$ and for HBLs, it is in the range $2.5\leq\delta\leq 3.0$. From detail analysis of several VHE spectra of HBLs it was observed that the flaring states can be in three categories, viz, (i) very high emission state with $2.5\leq\delta\leq 2.6$, (ii) high emission state with $3.02.6<\delta<$ 3.0 and (iii) low emission state with $\delta=3.0$ (Sahu et al., 2019). But for a tEHBL spectrum, it was shown that the single power-law behavior of the form $\Phi_{SSC}\propto E^{-\beta}_{\gamma}$ does not work. Thus, two power-laws for the background seed photons in the low energy tail region is assumed in the form (Sahu et al., 2020).
@@
$\Phi_{SSC}\propto\left\{\begin{array}[]{cr}E^{-\beta_{1}}_{\gamma},&\quad\mathrm{100\,GeV\,\lesssim E_{\gamma}\lesssim E^{c}_{\gamma}}\\ E^{-\beta_{2}}_{\gamma},&\quad\mathrm{E_{\gamma}\gtrsim E^{c}_{\gamma}}\\ \end{array}\right.,$ (4)
@@
with the spectral indices $\beta_{1}\neq\beta_{2}$ ($\beta_{i}>0$) and $E^{c}_{\gamma}$ is the cutoff energy scale around which the transition between zone-1 and zone-2 takes place. The value of $E^{c}_{\gamma}$ is different for different spectra. Thus, using Eq.(4), the observed spectrum for a temporary EHBL can be expressed as
@@
$F_{\gamma,obs}=e^{-\tau_{\gamma\gamma}}\times\begin{cases}F_{1}\,\left(\frac{E_{\gamma}}{TeV}\right)^{-\delta_{1}+3},&\quad\mathrm{100\,GeV\,\lesssim E_{\gamma}\lesssim E^{c}_{\gamma}}\,\,\,\,\,(\text{zone-1})\\ F_{2}\,\left(\frac{E_{\gamma}}{TeV}\right)^{-\delta_{2}+3},&\quad\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\mathrm{E_{\gamma}\gtrsim E^{c}_{\gamma}}\,\,\,\,\,(\text{zone-2})\end{cases}.$ (5)
@@
$F_{1}$ and $F_{2}$ are the normalization constants and $\delta_{i}=\alpha+\beta_{i}$ ($i=1,2$) are the free parameters in the photohadronic model that are to be adjusted by fitting to the observed VHE spectrum. Previously, the study of tEHBL-like behavior of Mrk421, Mrk501 and 1ES1959+650 in the context of two-zone photohadronic model have shown $2.5\leq\delta_{1}\leq 3.0$ and $3.0<\delta_{2}\leq 3.5$ (Sahu et al., 2020, 2021a, 2021b).

### v2

# Temporary EHBL-like behavior of Markarian 501 during July 2014 VHE flaring

@@ Abstract
Markarian 501, a BL Lac object well-known as a high energy gamma-ray source, has exhibited several epochs of very high energy (VHE) gamma-ray flaring events when its synchrotron peak frequency shifted above $10^{17}$ Hz, a signature of extreme behavior. From July 16 to July 31, 2014 such flaring events were observed for 15 days by various telescopes. On July 19 (MJD 56857.98), the X-ray outburst from the source was at its highest and on the same day an intriguing narrow peak-like feature around 3 TeV was observed by the MAGIC telescopes, a feature inconsistent with standard interpretations. Using the well-known two-zone photohadronic model, we study these VHE gamma-ray spectra on a day-by-day basis and offer explanation. Our two-zone photohadronic scenario shows that, on MJD 56857.98, the peak-like feature appears at a cutoff energy of $E^{c}_{\gamma}=3.18$ TeV. Below this energy the VHE spectrum increases slowly and is in high emission state. However, for $E^{c}_{\gamma}\,>3.18$ TeV, the spectrum falls faster, resulting in a mild peak-like feature, not prominent enough as claimed by the MAGIC collaboration.
journal: Eur. Phys. J. C
@@ 1 Introduction
Markarian 501 (Mrk 501) is a high-frequency peaked BL Lac object (HBL) at a redshift of $z=0.034$  Ulrich et al. (1975). It is the second well-known bright HBL after Markarian 421 (Mrk 421) and is one of the brightest extragalactic sources in the X-ray/TeV sky. After its discovery in very high energy (VHE) (>100>100 GeV) by the Whipple telescopes in 1996  Quinn and others (1996), several major outbursts in multi-TeV have been observed and have been the target of many multiwavelength campaigns mainly covering VHE flaring activities  Abdo and others (2011); Aleksić and others (2015a); Furniss and others (2015); Ahnen and others (2017, 2018). Mrk 501 is one of the few blazars which can be monitored even during the low emission periods. To understand different emission mechanisms and their characterizations, it is essential to have long-term and multiwavelength observations of the spectral energy distribution (SED) of the blazars. Also, simultaneous correlation studies in different wavelengths of the SED during the flaring can help us to constrain the emission mechanisms. Motivated by these goals, several multiwavelength and multi-instrument observation campaigns were organized. In 1997, from March to October, Mrk 501 had an unprecedented flare in VHE gamma-rays and the integral flux reached up to four times the magnitude of the Crab Nebula (CU) flux Petry and others (2000). Also, it was extremely active in the X-ray region during this period and the synchrotron peak had a dramatic shift to energies above 100 keV, which is more than two orders of magnitude compared to that of the typical non-flaring state.
@@
During May to July 2005, Major Atmospheric Gamma-Ray Imaging Cherenkov (MAGIC) telescopes made observations for 30 nights, and flaring above 100 GeV was observed with an order of magnitude flux variation Albert and others (2007). On the nights of June 30 and July 9, flux-doubling in the VHE band was observed within a time span of about 2 minutes. This is the fastest flux variation ever observed from Mrk 501 since its discovery. During this observation period, the synchrotron peak and the VHE peak were shifted towards higher energy limits. As a part of the long term multiwavelength campaign, Mrk 501 was observed by $\sim 30$ different instruments covering nearly the entire electromagnetic spectrum Ahnen and others (2017); Aliu and others (2016) from March 15 to August 1, 2009.
@@
Between March and July of 2012, another multiwavelength campaign was undertaken when more than 25 different telescopes, including MAGIC, and Very Energetic Radiation Imaging Telescope Array System (VERITAS) took part and observations at energies > 0.2>\,0.2 TeV were made Ahnen and others (2018). The highest peak flux was recorded on June 9, and most importantly, the X-ray spectrum and the VHE gamma-ray spectrum were observed to be extremely hard. During this period, it is observed that the synchrotron and the Synchrotron Self-Compton (SSC) peaks were above 5 keV and above 0.5 TeV respectively. The TeV Atmospheric Cherenkov Telescope with Imaging Camera (TACTIC) at Mt Abu, Rajastan, India, observed Mrk 501 from 15 April to 30 May, 2012 at energies above 850 GeV for 76.0 h. Between 22 to 27 May, 2012, TACTIC telescope observed relatively high VHE flux in the energy range between 850 GeV and 17.24 TeV Chandra and others (2017).
@@
Mrk 501 was monitored in the year 2014 with the First G-APD Cherenkov Telescope (FACT) Anderhub and others (2013); Biland and others (2014); Dorner and others (2015). Several high flaring states were observed. This triggered the High Energy Stereoscopic System (HESS) to undertake further observations. Enhanced VHE gamma-ray flux, comparable to the exceptional flare of 1997, was observed on the night of June 23-24, 2014 (MJD 56832) with the HESS phase-II telescopes. The energy of the VHE photons extended up to 20 TeV and the flux had a rapid variability. Estimates indicate that this is the highest flux ever observed from Mrk 501 with the HESS telescopes Abdalla and others (2019).
@@
Again, in July of 2014, flaring in X-ray and VHE gamma-ray were observed for about two weeks (from 16th to 31st) MAGIC Collaboration and others (2020b). The X-ray flaring was found to be exceptionally high, similar to the historic 1997 flaring. This amounts to the largest flux ever detected so far with the X-ray Telescope (XRT) on board the Neil Gehrels Swift Observatory Gehrels and others (2004) in its operational history. Also, the VHE gamma-ray emission above 0.15 TeV during this time was high, about 0.5 to 2 times of the CU flux. On July 19 (MJD 56857.98), the day of the highest X-ray activity, a narrow peak-like feature at $\sim 3$ TeV was observed by MAGIC telescopes MAGIC Collaboration and others (2020b).
@@
In the Universe, the most dominant sources of extragalactic gamma rays are the blazars Romero et al. (2017). Their SED is characterized by two non-thermal peaks Abdo and others (2010), and the most popular models interpret these peaks as the synchrotron peak is in the infrared to X-ray frequency range and the SSC peak is in the X-ray to gamma-ray range Fossati et al. (1998); Ghisellini et al. (1998); Urry (1999). The BL Lac objects belong to the blazar family and are classified according to the position of their synchrotron peak frequency $\nu^{p}_{s}$ in the SED; as low-frequency peaked (LBLs, $\nu^{p}_{s}<10^{14}\,\mathrm{Hz}$), intermediate-frequency peaked (IBLs, $10^{14}\,\mathrm{Hz}<\nu^{p}_{s}<10^{15}\,\mathrm{Hz}$), high-frequency peaked (HBLs, $10^{15}\,\mathrm{Hz}<\nu^{p}_{s}<10^{17}\,\mathrm{Hz}$ Abdo and others (2010); Padovani and Giommi (1995); Böttcher et al. (2013)) and extreme high-frequency peaked (EHBLs, $\nu^{p}_{s}>10^{17}\,\mathrm{Hz}$ Costamante and others (2001)) with extreme spectral properties.
@@
Low luminosity and limited variability of EHBLs hinder the detection of their gamma-rays by the current generation Imaging Atmospheric Cerenkov Telescopes (IACTs). Thus, only a handful of EHBLs have been detected so far and well-known among them are 1ES 0229+200, 1ES 0347-232, RGB J0710+591 and 1ES 1101-232 Costamante et al. (2018). It is observed that at least three nearby extensively studied HBLs, Mrk 421, Mrk 501 and 1ES 1959+650 have shown temporary EHBL (tEHBL) like behavior during several flaring epochs Sahu et al. (2020, 2021a, 2021b). During tEHBL-like behavior, the synchrotron peaks have shifted above $10^{17}$ Hz and the TeV spectra are harder. Also, due to the complex behavior of these sources with different spectral properties at VHE gamma rays, the possibility of different subclasses within the EHBL class cannot be ruled out Foffano et al. (2019).
@@
With the one-zone leptonic SSC model, it is difficult to explain the VHE spectrum of EHBL, since the leptonic model predicts a softer SSC spectrum in the Klein-Nishina regime, and is contrary to observations Paggi et al. (2009). One can circumvent this problem by adopting very large values of electron Lorentz factor, bulk Lorentz factor and a very low magnetic field Ahnen and others (2018); MAGIC Collaboration and others (2020a). Several alternative models such as the two-zone leptonic model, the inverse Compton scattering of relativistic electrons with the cosmic microwave background, the spine-layer structured jet model, and different hadronic models are proposed to explain these spectral behavior Ahnen and others (2018); Böttcher et al. (2008); MAGIC Collaboration and others (2019).
@@
VHE gamma rays from the extragalactic sources get attenuated by interacting with the extragalactic background light (EBL) and this attenuation is energy dependent. The shape of the observed spectrum changes at very high energies, so, EBL corrections to the observed spectrum have to be taken care of Franceschini et al. (2008); Domínguez and others (2011); Tavecchio (2014); Tavecchio and Bonnoli (2015). EHBLs have extreme spectral properties in TeV energy range, and probably these are ideal probes for EBL.
Mrk 501 had VHE flaring during 2005 and 2012 when the synchrotron peak had shifted temporarily above 1 keV and the SSC peak also shifted to higher energy, a clear signature of EHBL-like behavior. The shift of the synchrotron peak and the SSC peak to higher energies in the multiwavelength SED generally translate to a particularly hard X-ray and a very-high-energy gamma-ray spectra, difficult to explain using leptonic model.
Different temporal behaviour has been observed among the EHBLs. While the EHBLs, 1ES 0229+200 and 1ES 034-232 exhibit extreme properties constantly, the well-known HBLs, such as Mrk 421, Mrk 501 and 1ES 1959+650 enter to the EHBL family as temporary members, with peculiar spectral properties Ahnen and others (2018); Albert and others (2007); MAGIC Collaboration and others (2020a); Aleksić and others (2015b). The VHE spectrum of a tEHBL falls faster in the high energy regime (above a cutoff energy) as compared to low energy regime. This behavior is not observed in HBLs and EHBLs and can not be explained using the conventional photohadronic model. Thus, the photohadronic model was extended to two-zones, which could explain the VHE spectra extremely well Sahu et al. (2020, 2021a, 2021b). The VHE flaring event of July 2014 of Mrk 501 was very similar to the above flaring events. So, in this work, our goal is to explain the VHE flaring events observed during July 16 to July 31, 2014, and particularly, to interpret the unusual peak-like feature at about 3 TeV observed on July 19 (MJD 56857.98) in the context of the two-zone photohadronic model.
@@
The paper is organized as follows. In section 2, the two-zone photohadronic model is briefly discussed. In section 3, we discuss the flaring events of Mrk 501 during July 16 to July 31, 2014 (MJD 56854–56869). A detailed analysis of the VHE events observed by MAGIC for about two weeks is discussed in section 4 and a brief discussion is given in section 5.
@@ 2 Two-Zone Photohadronic Model
The synchrotron peak and the SSC peak of a BL Lac object are conventionally explained through the leptonic models. In these models, the region where the photons are emitted is a blob of comoving radius $R^{\prime}_{b}$ (where implies comoving frame), with a bulk Lorentz factor $\Gamma$ and a Doppler factor $\mathcal{D}$. For HBLs, $\Gamma\simeq\mathcal{D}$.
In the photohadronic scenario, accelerated protons are injected in the jet with a differential spectrum given as $dN/dE_{p}\propto E^{-\alpha}_{p}$, where the proton spectral index $\alpha\geq 2$ Dermer and Schlickeiser (1993). These protons interact with the background seed photons to produce $\Delta$-resonances, $p\gamma\rightarrow\Delta^{+}$. These background seed photons are produced from the leptonic processes (in the leptonic models). Decay of the $\Delta^{+}$ to neutral pion produces observed gamma-rays. The kinematical condition to produce $\Delta^{+}$ is given by Sahu (2019); Sahu et al. (2019),
$E_{p}\,\epsilon_{\gamma}=\frac{0.32\,\Gamma^{2}}{(1+z)^{2}}~\mathrm{GeV}^{2},$ (1)
where $E_{p}$ and $\epsilon_{\gamma}$ are respectively the proton energy and the seed photon energy in the observer’s reference frame. In Eq.(1), $z$ is the redshift of the source. The energy of the observed gamma-ray $E_{\gamma}$ from the $\pi^{0}$ decay carries 10% of the original proton energy.
It is observed that in the standard jet scenario of the BL Lac objects, the $p\gamma\rightarrow\Delta^{+}$ process is inefficient due to the low background photon density Pjanka et al. (2017).To avoid this problem and to explain the VHE gamma-ray events from HBLs, a double jet structure is assumed in the context of the photohadronic model Sahu (2019). In this model, it is assumed that during the VHE flaring of the source, an internal compact jet is formed within the larger jet Sahu et al. (2016, 2018b). The inner jet has radius $R^{\prime}_{f}$ and the outer jet has radius $R^{\prime}_{b}$ and $R^{\prime}_{f}<R^{\prime}_{b}$. Also, the photon density in the inner jet region, $n^{\prime}_{\gamma,f}$, is larger than the photon density in the outer jet region, $n^{\prime}_{\gamma}$ ($n^{\prime}_{\gamma,f}\,\gg\,n^{\prime}_{\gamma}$). During the VHE flaring, the accelerated high energy protons interact with the seed photons in the inner jet region to produce $\Delta$-resonance. Due to high density photons in the inner jet region, the efficiency of the production of $\Delta$-resonance is relatively high as compared to the standard jet scenario and this avoids the necessity for super-Eddington luminosity of protons.
Since $R^{\prime}_{f}<R^{\prime}_{b}$, the photon density in the inner jet region $n^{\prime}_{\gamma,f}$, can not be observed directly. But, in the outer region the photon density $n^{\prime}_{\gamma}$, can be calculated from the observed flux. Also, due to the adiabatic expansion of the inner jet, its photon density decreases when it crosses into the outer jet. To connect the photon densities in the inner and outer regions, we assume a scaling relation given by Sahu et al. (2019, 2013)
@@
From the previous analysis of the VHE spectra of dozens of HBLs in the photohadronic model, we found that to produce the observed $E_{\gamma}$, the background seed photon energy $\epsilon_{\gamma}$ always lies in the low energy tail region of the SSC spectrum which satisfies Eq. (1), and the SSC flux is a perfect power-law, given by $\Phi_{SSC}\propto\epsilon^{\beta}_{\gamma}\propto E^{-\beta}_{\gamma}$, with $\beta$ the photon spectral index Sahu (2019). However, the low energy tail region of the SSC spectrum is model dependent as there is no observation in this energy band. The observed GeV-TeV photon flux from a HBL is given as Sahu (2019); Sahu et al. (2019)
$F_{\gamma,\mathrm{obs}}(E_{\gamma})$ $=F_{0}\left(\frac{E_{\gamma}}{\mathrm{TeV}}\right)^{-\delta+3}e^{-\tau_{\gamma\gamma}(E_{\gamma},z)}$
$=F_{\gamma,\mathrm{in}}(E_{\gamma})\,e^{-\tau_{\gamma\gamma}(E_{\gamma},z)}.$ (3)
@@
In the above equation, $F_{0}$ is the flux normalization, $F_{\gamma,in}$ is the intrinsic VHE gamma-ray flux and the exponential factor corresponds to the survival probability of VHE photon due to its interaction with the EBL photons which has the optical depth $\tau_{\gamma\gamma}$. For the EBL correction to the VHE spectrum, here, we use the EBL model of Franceschini et al Franceschini et al. (2008). The exponent $\delta=\alpha+\beta$ and for HBLs, it is in the range $2.5\leq\delta\leq 3.0$. From detailed analysis of several VHE spectra of HBLs it was observed that the flaring states can be in three categories, viz, (i) very high emission state with $2.5\leq\delta\leq 2.6$, (ii) high emission state with $2.6<\delta<3.0$ and (iii) low emission state with $\delta=3.0$ Sahu et al. (2019). But for a tEHBL spectrum, it was shown that the single power-law behavior of the form $\Phi_{SSC}\propto E^{-\beta}_{\gamma}$ does not workSahu et al. (2020, 2021a, 2021b). Thus, two power-laws for the background seed photons in the low energy tail region is assumed in the form Sahu et al. (2020),
@@
$\Phi_{SSC}\propto\left\{\begin{array}[]{cr}E^{-\beta_{1}}_{\gamma},&\quad\mathrm{100\,GeV\,}\lesssim E_{\gamma}\lesssim E^{c}_{\gamma}\\ E^{-\beta_{2}}_{\gamma},&\quad E_{\gamma}\gtrsim E^{c}_{\gamma}\\ \end{array}\right$., (4)
@@
with the spectral indices $\beta_{1}\neq\beta_{2}$ ($\beta_{1},\beta_{2}>0$) and $E^{c}_{\gamma}$ is the cutoff energy scale around which the transition between zone-1 and zone-2 takes place. The value of $E^{c}_{\gamma}$ is different for different spectra. Thus, using Eq.(4), the observed spectrum for a tEHBL can be expressed as
@@
$F_{\gamma,obs}=e^{-\tau_{\gamma\gamma}}\begin{cases}F_{1}\left(\frac{E_{\gamma}}{TeV}\right)^{-\delta_{1}+3},&\,(\text{zone-1})\\ F_{2}\left(\frac{E_{\gamma}}{TeV}\right)^{-\delta_{2}+3},&\,\,(\text{zone-2})\end{cases},$ (5)
@@
were $F_{1}$ and $F_{2}$ are the normalization constants and $\delta_{i}=\alpha+\beta_{i}$ ($i=1,2$) are the free parameters in the photohadronic model that are to be adjusted by fitting to the observed VHE spectrum. In Eq.(5), the zone-1 corresponds to the energy range $\mathrm{100\,GeV\,\lesssim E_{\gamma}\lesssim E^{c}_{\gamma}}$ and the zone-2 corresponds to $\mathrm{E_{\gamma}\gtrsim E^{c}_{\gamma}}$. Previously, the study of tEHBL-like behavior of Mrk 421, Mrk 501 and 1ES 1959+650 in the context of two-zone photohadronic model has shown $2.5\leq\delta_{1}\leq 3.0$ and $3.0<\delta_{2}\leq 3.5$ Sahu et al. (2020, 2021a, 2021b).